# The impact of gravitational waveform model systematics on the measurement of the Hubble Constant

### v1

@@ Acknowledgements.
N.K., P.T.H.P., T.D. acknowledge support from the Daimler and Benz Foundation for the project “NUMANJI”. T.D. acknowledges support from the European Union (ERC, SMArt, 101076369). Views and opinions expressed are those of the authors only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them. P.T.H.P. is supported by the research program of the Netherlands Organisation for Scientific Research (NWO). We acknowledge usage of computer time on Lise/Emmy of the North German Supercomputing Alliance (HLRN) [project bbp00049], on HAWK at the High-Performance Computing Center Stuttgart (HLRS) [project GWanalysis 44189], and on SuperMUC NG of the Leibniz Supercomputing Centre (LRZ) [project pn29ba].

### v2

@@ Acknowledgements.
N.K., P.T.H.P., and T.D. acknowledge support from the Daimler and Benz Foundation for the project “NUMANJI”. T.D. acknowledges support from the European Union (ERC, SMArt, 101076369). Views and opinions expressed are those of the authors only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them. P.T.H.P. is supported by the research program of the Netherlands Organisation for Scientific Research (NWO). We acknowledge usage of computer time on Lise/Emmy of the North German Supercomputing Alliance (HLRN) (Project No. bbp00049), on HAWK at the High-Performance Computing Center Stuttgart (HLRS) (Project No. GWanalysis 44189), and on SuperMUC NG of the Leibniz Supercomputing Centre (LRZ) (Project No. pn29ba).